OpenStudy (anonymous):

8 cos 2x intregrate it

OpenStudy (saifoo.khan):

u may find it on xtreme papers

OpenStudy (watchmath):

\(4\sin(2x)\)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\](8cos 2x=(8/2) sin 2x=4sin 2x

OpenStudy (anonymous):

WHo cares

OpenStudy (saifoo.khan):

ariba does.

OpenStudy (anonymous):

So

OpenStudy (saifoo.khan):

so, she just care for him.